OpenStudy (anonymous):

please help me in solving this problem of 3D??

OpenStudy (anonymous):

find the distance of the point (1,-2,3) from the plane x-y+z=5 measured parallel to the line x/2=y/3=z/-6

OpenStudy (amistre64):

the normal of the plane is: <1,-1,1> apply this to the point: (1,-2,3) x = 1+t y = -2-t z = 3+t any point on our plane satisfies: x-y+z=5, so lets sub this in to find "t" along the normal 1+t-(-2-t)+3+t = 5 1+t+2+t+3+t = 5 6+3t = 5 t = -1/3; this should be the distance the point is hovering away from the plane itself, a perp measure.

OpenStudy (amistre64):

x = 1 -1/3 = 2/3 y = -2+1/3 = -5/3 z = 3- 1/3 = 8/3 the point on the plane directly beneath directly beneath the given should be: (2/3 ,-5/3, 8/3); which turns out good. the distance from this point to our hovering point is then: 3/3, -6/3, 9/3 -(2/3 ,-5/3, 8/3) --------------- 1/3, 11/3, 17/3 sqrt(1/9 + 121/9 + 289/9) sqrt(411)/3 above the plane; but thats just side track stuff. we can use the same concept to find the distance along the other line given

OpenStudy (amistre64):

x/2=y/3=z/-6 t=x/2 t=y/3 t=z/-6 2t+0= x 3t+0= y -6t+0=z we have the vector part, now apply our given point 2t+1= x 3t-2= y -6t+3=z and find where it hits the plane 2t+1 -3t+2 -6t+3 = 5 -7t+6 = 5 t = 1/7 2(1/7)+1 = x = 9/7 3(1/7)-2 = y = -11/7 -6(1/7)+3=z = 15/7 now we just need to measure the length between the given point and the plane point by the vector between them

OpenStudy (amistre64):

7/7, -14/7, 21/7 -(9/7,-11/7, 15/7) ----------------- <-2/7, -3/7, 6/7> sqrt(4+9+36)/7 sqrt(49)/7 = 1 with any luck, that works :)